# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

### v2

@@ 4.2 Results and Analysis
Monolingual Cosslingual
Task Lang. InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa
1. Polarization Amharic 77.96 78.92 81.93 80.64 53.71 74.98 23.26 49.44 5.63 0.00 0.00 83.21
Arabic 58.39 61.94 67.19 70.42 52.48 54.68 29.68 35.87 27.76 2.78 37.79 25.63
English 72.09 69.84 75.43 76.08 72.77 74.94 1.90 51.68 53.41 46.89 33.44 53.45
German 29.75 63.46 67.50 67.78 58.45 64.38 0.62 51.97 64.59 41.87 46.05 25.94
Hausa 59.57 66.41 67.43 66.41 59.32 65.62 8.21 22.92 6.47 1.50 6.25 25.13
Spanish 38.69 64.04 70.98 57.00 59.31 54.16 45.70 68.66 67.26 21.09 68.34 62.17
Urdu 1.07 66.35 79.95 43.91 65.63 60.68 1.60 30.36 0.00 0.00 3.68 68.67
2. Types Amharic 24.22 38.13 43.65 25.71 17.93 25.00 11.56 20.65 2.49 0.00 0.00 26.76
Arabic 22.97 40.23 42.12 37.58 27.53 35.22 11.49 23.73 8.78 2.11 12.47 7.93
English 16.04 23.25 31.38 24.10 21.07 17.70 10.84 19.69 16.42 12.22 11.70 3.93
German 13.52 58.58 61.19 58.56 54.13 40.64 12.03 35.86 29.59 9.72 23.98 11.88
Hausa 18.56 19.14 17.38 18.09 19.61 18.87 3.20 9.97 3.91 1.39 5.78 5.90
Spanish 43.26 66.07 67.76 58.07 57.94 43.40 7.89 47.06 15.38 0.43 32.02 14.60
Urdu 27.14 51.94 51.60 45.38 38.02 33.13 3.77 13.62 6.33 0.00 3.94 20.30
3. Manifestations Amharic 43.52 47.56 47.63 43.17 33.18 43.29 15.57 27.07 8.64 0.00 0.00 43.58
Arabic 40.05 51.55 52.52 55.61 42.18 47.73 16.56 30.68 22.14 0.00 19.57 17.24
English 14.40 15.01 19.39 18.61 18.60 15.15 7.16 10.16 10.05 5.62 10.69 8.85
German 38.49 49.88 52.74 51.70 46.85 51.91 2.38 36.12 27.05 0.00 23.38 12.93
Hausa 19.23 20.04 19.18 18.89 18.74 18.93 5.74 5.86 3.77 3.12 6.19 6.43
Spanish 38.94 50.00 51.04 45.02 45.17 35.09 2.34 40.63 11.83 0.40 35.99 23.56
Urdu 34.26 52.20 53.64 41.32 45.90 47.01 2.10 19.16 11.54 0.00 1.88 48.58
Table 3: Average F1-Macro for all three tasks. In the monolingual settings, we train and evaluate the model for each language separately. In the crosslingual setting, we train on all languages within a language family (AfroAsian: Amharic, Arabic, Hausa, and Urdu), Germanic: English, German, and Spanish) except the target language, and evaluate on the test set of the target language. The best performance scores are highlighted in blue and orange, respectively.
4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
For Task 1, binary polarization classification, RemBERT consistently achieved the highest macro F1-scores for most languages, while XLM-R performed best in Arabic, English, and German. In the multi-label polarization types classification (Task 2), RemBERT again provided the top scores for Spanish, except for Hausa and Urdu, where mBERT and LaBSE performed best, respectively. For Task 3, multi-label polarization manifestations classification, the results were generally lower across all models, with RemBERT delivering the highest scores for most languages, except for Arabic and Hausa.
4.2.2 Cross-Lingual MLM Results
For cross-lingual transfer experiments, each MLM was trained on all other languages within the same language family (Afro-Asian or Germanic), excluding the target language, and evaluated on the target language’s test set for the same three tasks.
In the cross-lingual setup for polarization classification (Task 1), mDeBERTa achieved the highest F1-macro for Amharic, English, Hausa, and Urdu. For Arabic, German, and Spanish, mBERT, RemBERT, and LaBSE performed slightly better than the other models, respectively.
For Task 2, polarization types classification, performance was generally lower as expected in the cross-lingual scenario. LaBSE yielded the best results for four languages, while mDeBERTa performed better for Amharic and Urdu.
Similarly, for Task 3, polarization manifestation classification, the results were quite low overall, but LaBSE and mDeBERTa showed relatively better performance compared to the other models.
This suggests that general polarization detection transfers more easily across languages than the more nuanced identification of polarization types and manifestations, which are culturally and linguistically specific. Effective cross-lingual transfer likely requires aligned annotation schemes and fine-tuning strategies sensitive to these differences.
4.2.3 Zero and Few-shot LLM Performance
We evaluated seven state-of-the-art LLMs (Qwen3-8B, LLaMA-3.1-8B, Mixtral-8x7B, GPT-4o, GPT-4o-mini, GPT-4.1-mini, and Llama-4-Scout-17B-16E-Instruct) across languages. The prompts used for evaluation are provided in the Appendix B. We explicitly provided definitions of tasks and leveraged the models’ chain-of-thoughts capabilities by encouraging reasoning before labeling. In the few-shot setting, we provided 4 example instances for each task. To assess the impact of prompt language, we also compared model performance when prompted in English versus the respective target language.
As Table  4 illustrates, the closed-source models GPT-4o, GPT-4o mini, and GPT-4.1 mini, consistently outperform the other four LLMs, demonstrating superior performance in polarization recognition. However, their performance on Hausa remains notably low, highlighting both the challenge of understanding low-resource languages and the value of our multilingual dataset in exposing the limitations.
In contrast, smaller open-source models, such as Qwen3-8B, LLaMA-3.1-8B, and Mixtral-8x7B, achieve lower F1-Macro scores and perform inconsistently across languages. Their F1-Macro scores improve by 5 10% when provided with few-shot examples. However, the larger LLaMA model, LLaMA-4-scout-17B, demonstrates performance comparable to GPT-4o in polarization classification.
An interesting phenomenon is observed: Smaller models show improved classification performance in the few-shot setting, while larger models such as the GPT series and LLaMA-4-17B do not consistently gain additional improvement with few-shot examples. This suggests that few-shot learning offers advantages to smaller models than to larger ones.
Task Lang GPT 4o GPT 4o-mini GPT 4.1-mini Mistral-7B Qwen3-8B Llama-3.1-8B Llama4-17B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot
1. Polarization Amharic 71.30 68.74 55.40 58.04 63.52 63.20 44.84 48.24 55.60 65.51 50.98 58.84 69.96 67.91
Arabic 71.51 79.65 58.21 68.41 76.20 79.58 33.65 53.06 65.95 70.76 61.51 67.21 73.58 77.45
English 75.46 79.56 64.78 80.30 77.43 77.48 53.78 70.00 69.09 66.86 66.55 69.90 76.90 75.48
German 72.04 71.12 63.03 68.49 73.77 69.60 56.17 61.77 65.42 68.76 59.90 65.40 70.89 70.35
Hausa 34.73 44.87 28.21 34.07 39.45 48.06 19.82 28.15 45.91 49.18 38.73 40.80 47.85 54.53
Spanish 72.69 65.76 62.51 67.96 71.28 69.22 51.24 61.00 65.42 68.76 60.20 62.53 66.11 66.27
Urdu 72.25 78.83 59.19 69.92 73.15 72.63 48.27 54.79 66.99 70.89 55.61 59.77 73.93 68.50
2. Types Amharic 39.67 36.59 33.8 34.89 32.49 34.44 44.21 63.24 63.07 75.90 78.26
Arabic 35.00 39.30 37.18 36.82 33.07 36.77 49.39 53.91 53.22 52.86 55.03
English 34.96 35.67 33.79 38.73 35.22 38.22 48.28 68.50 67.37 62.71 67.09
German 58.86 58.58 54.36 57.94 52.70 55.15 44.10 71.82 70.62 73.89 76.20
Hausa 18.96 19.11 16.90 18.15 17.00 18.86 43.34 67.24 70.47 71.75 74.66
Spanish 48.98 52.93 52.82 55.09 48.89 49.76 43.15 65.96 67.87 69.82 71.43
Urdu 55.6 59.64 53.97 54.78 51.82 55.09 48.40 63.36 64.83 60.01 61.52
3. Manifestations Amharic 44.09 45.59 40.01 44.78 40.53 42.49 0 0 0 0 0
Arabic 51.90 56.90 46.96 51.01 52.57 60.76 0 0 0 0 0
English 22.95 25.87 20.43 25.78 28.41 26.67 0 0 0 0 0
German 42.84 45.16 43.33 53.78 37.62 41.23 0 0 0 0 0
Hausa 8.58 8.14 8.42 8.37 8.55 9.33 0 0 0 0 0
Spanish 41.00 40.99 39.85 46.33 38.82 40.66 0 0 0 0 0
Urdu 52.89 54.43 48.25 58.17 48.34 51.89 0 0 0 0 0
Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
@@ 5 Discussion
Our multilingual and cross-cultural investigation into online polarization reveals that polarization is a deeply contextual and event-driven phenomenon. Its prevalence and intensity are closely tied to the sociopolitical climate, public discourse, and digital environment of each region. In societies experiencing major crises, such as the war in Ethiopia, religious tensions in Pakistan, and international conflicts like the wars in Ukraine and Gaza that influence the US election, we find polarization to be especially amplified on online platforms. Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
The types of polarization evident in online discourse are shaped by the most salient local issues: political (e.g., in Ethiopia, the US, and Germany), religious (e.g., in Pakistan), ethnic (e.g., in Ethiopia), or gender-based (e.g., in Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict. This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
Polarization also manifests in varied rhetorical forms, including stereotyping/deindividuation, vilification, and exclusionary or hostile language. These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
From a computational perspective, our results demonstrate both the feasibility and the challenges of detecting polarization in multilingual settings. Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec. 4). However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding. Performance on these more nuanced tasks remains limited. This highlights important directions for future work, including the integration of cultural signals and context or event understanding, the development of more robust multilingual embeddings, and the pursuit of consistent annotation (with a great deal of annotator training and follow up, specially for the crowdsourcing annotations). The wide variability across languages, contexts, and events further affirms the necessity of designing language, culture, and event specific approaches, particularly for underrepresented regions in global NLP research.
Taken together, these findings highlight that effective detection and mitigation of online polarization must move beyond generic or monocultural solutions. There is a pressing need for context-aware, adaptable methodologies that acknowledge both the universal and the local characteristics of polarized discourse, ensuring relevance and effectiveness across the world’s diverse digital landscapes.
@@ 6 Conclusion
In this study, we introduced POLAR, the first multilingual, multicultural, and multievent dataset for benchmarking online polarization. Our findings showed that polarization was deeply context-dependent, manifesting in diverse types and rhetorical forms shaped by local sociopolitical dynamics. Theoretically, we provide new evidence that polarization is a multifaceted, cross-lingual phenomenon strongly influenced by language, culture, and current events, rather than being uniform across settings. On a practical level, our benchmark revealed that while binary detection of polarization was feasible with current models, accurately identifying specific types and manifestations remained a persistent challenge, especially for multilingual and low-resource contexts.
This work established a foundation for further computational studies of digital polarization and for developing more culturally robust moderation and intervention tools. Looking ahead, future research should integrate richer contextual and cultural signals into model architectures, refine annotation guidelines for cross-regional consistency, and expand evaluations to include additional languages, social domains, and further events. By releasing our data and benchmarks, we aimed to catalyze further innovation toward nuanced detection and effective mitigation of online polarization worldwide.
@@
As Table 5 demonstrated, the inter-annotation agreement scores range from 0.2 to 0.6, corresponding to fair to moderate levels of agreement according to established benchmarks (landis1977kappa). This level of agreement further highlights the inherent difficulty of polarization recognition and underscores the necessity and value of our proposed work.
Language Annotation Platform Kappa Score
Amharic Potato 0.65
Arabic Potato 0.296
English MTurk 0.52
German Prolific 0.29
Hausa Potato 0.48
Spanish MTurk 0.24
Urdu Potato 0.55
Table 5: Dataset annotation platforms and inter-annotator agreement scores
Appendix A Annotation Guidelines

### v3

@@ 4.2 Results and Analysis
SLMs Models: Table 2 presents the results of six small language models. Overall, RemBERT and LaBSE show comparable performance across most languages, achieving the best or the second best macro-F1 scores. RemBERT is designed to balance representation across high-, mid-, and low-resource languages, while LaBSE relies on bilingual sentence-level alignment between English and low-resource languages. These training strategies enhance their ability to mid- and low-resource language understanding, which is reflected in their consistent improvements over mBERT and XLM-R, particularly for languages such as Amharic, Odia, Italian.
The twitter-roberta-hate model is finetuned on English Twitter dataset covering emoji, stance, hate speech, and emotion. This training boosts its performance on English polarization detection, polarization types classification and polarization manifestation recognition. AfroXLMR-large-76L is finetuned upon XLM-R on African languages datasets. As a result, it performs well on African language in our dataset, including Arabic, Hausa, and Swahili.
Polarization detection is comparatively easier for multilingual BERT-based models, which achieve relatively high macro-F1 scores. In contrast, recognizing polarization types (politics, gender, racial, religious, and others) is substantially more challenging. Performance drops even further for polarization manifestation recognition (stereotype, vilification, dehumanization, extreme language, lack of empathy, invalidation), where macro-F1 scores decrease markedly. This gap highlights the limitations of current models in capturing fine-grained and implicit polarization manifestations and the importance of POLAR  for advancing research on nuanced polarization understanding.
LLMs Performance: The Tables 2, 3, and 4 present an overall picture of model performance in polarization detection, types, and manifestations across the languages, using the Macro-F1 metric as a measure of accuracy. Across all experiment setups, models consistently achieve their highest scores in Polarization Detection, followed by the classification of Polarization Types, while identifying Polarization Manifestations remains the most challenging task for both encoder models and LLMs likely due to the latent nature of how polarization is expressed, requiring semantic understanding that state-of-the-art models find difficult to generalize across diverse cultural contexts.
Table 2 presents the the performance of encoders (mBERT, XLM-R, RemBERT, LaBSE, twitter-roberta-hate, and afro-xlmr-large), demonstrating significant stability across languages, especially in the PolarDetect task. On the other hand, Table 3 presents a more mixed picture through zero-shot LLM experiments. While high-resource languages like English (eng) and Chinese (zho) maintain high detection scores (reaching 79.84% and 82.48% respectively), the models exhibit a significant drop in performance for languages with unique scripts or less resources. For example, Khmer (khm) detection scores drop to a range between 8.65% and 13.32%, indicating that without prior examples, these generative models may not generalize. Table 4 demonstrates the transformative impact of few-shot prompting, where providing the LLMs with a few examples significantly improves performance for many languages. For instance, Urdu (urd) has the detection score rise from 40.97% in the zero-shot setting to 73.81% with GPT4.1 in the few-shot setting. In this few-shot setting, LLMs also outperform BERT-family encoders in the complex PolarManifest task, for example, with Arabic (arb) reaching a peak of 75.49%. In summary, while BERT-family models remain the most efficient for binary detection, LLMs with few-shot prompting show potential for handling more complex, fine-grained classification tasks when sufficient examples are provided.
Finally, some of the linguistic families exhibit clustered behaviors. For instance, a language cluster emerges among South Asian languages like Bengali (ben), Hindi (hin), Nepali (nep), and Oriya (ori), as these languages exhibit similar performance improvements in the few-shot LLM experiments, compared to zero-shot settings (Table 4). Nevertheless, the most dominant predictor of model performance seems to be resource-tier of the language, which often aligns with geographic and economic regions.
# Lang. PolarDetect PolarType PolarManifest
Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS
1 eng 77.98 77.79 74.90 75.82 79.82 76.81 79.74 72.17 73.52 70.48 77.33 76.96 74.01 79.24 62.37 63.40 61.56 63.41 64.35 63.77 69.40
deu 65.49 65.79 65.93 64.34 71.26 71.66 69.61 65.03 67.75 68.37 69.58 72.44 68.31 69.61 58.36 60.76 55.97 60.59 53.72 57.59 65.03
urd 57.89 55.42 55.48 43.96 62.62 73.81 66.26 40.80 49.04 47.02 45.40 51.85 48.14 49.39 43.73 57.34 54.55 59.44 58.42 49.75 56.67
ben 66.84 63.00 59.91 65.61 79.64 78.89 78.32 70.14 72.51 71.14 73.37 75.11 71.12 76.71 57.01 53.29 51.41 53.12 53.72 58.90 60.36
hin 52.92 43.45 41.42 52.13 58.54 62.61 56.26 59.85 54.86 55.54 69.68 71.31 66.72 70.99 47.77 51.70 54.76 59.41 58.65 54.29 57.49
ori 60.85 57.23 60.81 43.60 63.62 65.71 66.09 63.98 66.26 67.26 50.38 75.68 67.56 73.38 54.71 56.56 53.24 48.69 59.73 57.57 62.57
nep 64.38 74.27 64.25 69.50 84.64 82.44 81.20 68.49 73.98 71.72 77.16 81.57 74.45 82.38 59.81 64.29 60.89 68.35 70.57 72.03 76.92
pan 51.19 60.47 56.38 52.95 69.77 72.07 69.76 61.04 64.38 62.86 64.38 69.43 66.76 70.57 50.94 53.07 53.95 55.51 57.22 58.66 62.82
spa 68.01 70.43 67.35 66.91 72.15 74.05 74.61 68.46 68.96 64.16 70.36 71.67 68.65 73.39 59.49 58.69 58.10 59.80 60.00 62.54 66.29
ita 59.37 55.83 48.57 51.98 65.92 71.82 72.02 60.67 63.55 63.15 64.49 70.46 68.51 72.98
rus 74.96 73.70 62.74 70.59 74.32 74.04 77.45 69.78 70.74 63.57 74.10 73.97 71.19 78.71
pol 58.07 62.86 53.14 59.10 68.38 71.54 79.17 70.56 66.92 69.71 71.19 76.91 72.50 79.79
fas 40.77 41.32 40.45 39.32 54.58 54.22 50.53 57.14 59.51 66.76 59.26 69.72 62.67 65.54 58.08 55.41 54.25 56.35 57.53 59.24 62.47
2 hau 55.01 53.92 55.60 50.47 56.24 55.08 59.28 57.89 55.73 55.49 54.61 57.79 56.69 60.64 50.80 49.01 47.69 46.19 45.98 50.87 49.94
arb 75.54 72.88 73.42 72.21 79.58 78.37 78.40 64.36 65.53 66.16 70.07 71.44 70.50 72.89 64.89 66.55 61.96 67.07 67.6 67.58 75.49
amh 38.51 35.70 37.66 28.59 74.58 51.81 54.67 50.41 59.07 64.44 55.74 76.65 59.64 66.07 50.90 55.80 51.50 53.36 58.72 52.91 61.37
3 zho 56.69 68.17 56.79 66.71 69.81 78.22 74.57 71.13 78.00 72.03 76.32 76.87 79.12 77.00 69.97 67.39 61.25 62.78 63.98 72.01 76.47
mya 62.17 71.12 60.51 50.96 73.90 68.46 63.07 56.23 63.35 68.11 58.24 71.63 65.54 63.89
4 khm 24.70 13.66 15.75 12.98 18.90 16.52 15.69 53.99 51.21 62.92 51.81 57.24 51.99 51.41 49.45 50.55 49.55 52.19 50.47 48.27 50.11
5 swa 53.67 58.43 58.75 59.20 67.70 64.56 66.61 61.14 65.17 65.74 66.71 70.80 67.59 69.81 56.27 54.88 53.13 53.58 56.04 57.62 61.13
6 tel 58.52 56.59 53.40 54.56 60.02 59.74 52.15 52.15 55.13 56.16 56.64 57.24 56.24 53.92 53.79 55.79 55.41 54.61 59.03 56.71 52.53
7 tur 72.29 68.12 66.23 68.70 79.62 76.45 78.28 64.02 66.43 62.74 71.08 76.03 69.09 77.67 60.74 58.18 60.60 64.70 64.83 64.14 72.87
Table 4: F1-Macro scores(in %) from the few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
@@ 5 Discussion
Performance within the same language family:
It is not necessary that models uniformly achieve a high performance across all languages within the same language family. For example, while a model may perform strongly on Chinese, its performance on Burmese still lags behind, demonstrating that high-resource language performance does not guarantee transferability to related lower-resource languages such as Burmese.
Few-shot vs. Zero-shot:
Few-shot performance is not always superior to zero-shot results. For larger models, such as Gemma3 and GPT-OSS, the few-shot setting underperforms zero-shot, indicating that large LLMs are inherently capable of recognizing polarized sentences without requiring in-context exemplars. In contrast, for smaller models, including Qwen2.5, Qwen3, LLaMA3.1, and Mistral3, in-context examples (3-shot prompting) can improve performance, showing that smaller LLMs benefit more from in-context learning.
LLMs vs. SLMs:
SLMs demonstrate improved performance in detecting polarization after fine-tuning, but still struggle to distinguish polarization types and manifestations, reflecting limited semantic knowledge of domain-specific concepts. Conversely, LLMs generally exhibit a stronger understanding of social-science polarization constructs, enabling better recognition of polarization types and manifestations, yet they remain weaker in direct polarization classification tasks. This suggests that LLMs possess richer implicit knowledge of social-science constructs, whereas SLMs currently lack comparable semantic grounding for fine-grained polarization distinctions.
6 Error Analysis: Misclassification Cases
Based on our error analysis (See detail in Appendix G Table 8), we identified a misalignment between the model’s classification logic and human judgment of polarization. The model relies on a deterministic, surface-level heuristic: it classifies text as polarized only when it detects explicitly named and opposed groups within the sentence (e.g., "Israeli forces vs. Palestinians"). Conversely, if the text expresses hostility but names only one group or relies on implied opposition, the model defaults to labeling it as non-polarized.This error stems from models’ reliance on textual pattern alone, while human annotators draw upon cultural and contextual knowledge to interpret hostility and implicit group conflict.
7 Implications
Theoretical Implications: Our dataset highlights the complexity of online polarization, emphasizing its deep cultural and contextual nature. It reveals the current limits of NLP models in detecting implicit rhetorical tactics, underscoring the need for culturally-aware frameworks.
Practical Implications: The dataset provides a valuable benchmark for developing and evaluating models capable of detecting nuanced forms of online polarization across multiple languages and contexts. It supports the creation of more culturally sensitive and robust tools for the monitoring and mitigation efforts of online discourse.
Methodological Implications: Our multi-label, multi-platform annotation approach underscores the importance of culturally sensitive and detailed labeling strategies. The variability of model performance across languages and contexts indicates a pressing need for methods that integrate cultural signals, multimodal data, and contextual embeddings to improve robustness and reduce performance disparities in social NLP applications.
@@ 8 Conclusion
In this study, we introduced POLAR, a comprehensive, multilingual, and multi-event dataset designed to advance the understanding and detection of online polarization across diverse linguistic and cultural contexts. By annotating over 110,000 instances along three critical dimensions, we created a nuanced, fine-grained resource. This dataset captures the complex rhetorical tactics and social dimensions that underpin polarized discourse. Our extensive benchmarking of SOTA SLMs and LLMs reveals that while current models are reasonably effective at binary polarization detection, they face significant challenges in accurately identifying polarization types and rhetorical manifestations, especially in low-resource and culturally nuanced settings. These findings emphasize the deep contextual and implicit nature of online polarization and highlight the limitations of existing NLP approaches. Importantly, our work underscores the critical need for culturally aware, adaptable, and context-sensitive models to effectively monitor and mitigate digital polarization globally. The resources and benchmarks provided herein aim to catalyze future research, fostering the development of more inclusive and robust tools for analyzing social phenomena in multilingual and multicultural online environments.
@@
Appendix A Language and its language family
The POLAR dataset covers 22 languages from seven linguistic families. The coverage language and their corresponding branches is presented below:
# Lang. ISO-639 Language Family Sub-branch
1 English eng Indo-European Germanic
German deu Indo-European Germanic
Urdu urd Indo-European Indo-Aryan
Bengali ben Indo-European Indo-Aryan
Hindi hin Indo-European Indo-Aryan
Odia ori Indo-European Indo-Aryan
Nepali nep Indo-European Indo-Aryan
Punjabi pan Indo-European Indo-Aryan
Spanish spa Indo-European Romance
Italian ita Indo-European Romance
Russian rus Indo-European Slavic
Polish pol Indo-European Slavic
Persian fas Indo-European Iranian
2 Hausa hau Afro-Asiatic Chadic
Arabic arb Afro-Asiatic Semitic
Amharic amh Afro-Asiatic Semitic
3 Chinese zho Sino-Tibetan Sinitic
Burmese mya Sino-Tibetan Tibeto-Burman
4 Khmer khm Austroasiatic Mon-Khmer
5 Swahili swa Niger–Congo Bantu
6 Telugu tel Dravidian Dravidian
7 Turkish tur Turkic Turkic
Table 5: Language covered and its language families
Appendix B Data Statistics
Table 6 summarizes the distribution of instances across data sources in the POLAR dataset. The data primarily originate from major social media platforms, with Twitter contributing over half of the instances. Additional data are drawn from news websites, online forums, and other social platforms to ensure coverage of diverse discourse contexts. A smaller portion of the data comes from existing datasets that were re-annotated to align with our polarization framework.
Source No. of Instances %
X (Twitter) 58,984 53.3%
News/websites 15,208 13.7%
Bluesky 8,099 7.3%
YouTube 7,159 6.5%
Reddit 5,705 5.2%
Existing Dataset 4,794 4.3%
Weibo 4,550 4.1%
Facebook 2,928 2.6%
Zhihu 1,088 1.0%
Threads 852 0.8%
Tieba 783 0.7%
Wikipedia 500 0.5%
Total 110,650 100%
Table 6: Data sources of the POLAR dataset.
B.1 Dataset Composition by Language
Table 7 presents a language-wise overview of the POLAR dataset, including data sources, targeted events or topics, and inter-annotator agreement. For each language, data collection was tailored to platform availability and sociopolitical relevance, resulting in variation in event focus and discourse type. Inter-annotator agreement is reported primarily using Fleiss’s kappa, with alternative reliability measures noted where applicable.
B.2 Polarization Statistics
Figure 3 shows the distribution of polarization types, and Figure 4 illustrates the distribution of polarization manifestations across languages.
Figure 3: PolarType  by Languages, for each language, the numeric value shows the count of instances assigned to a given polarization type, while the percentage reflects its share of the total annotated instances for that language.
Figure 4: PolarManifest-by Languages. For each language, the numeric value shows the count of instances assigned to a given polarization manifestation, while the percentage reflects its share of the total annotated instances for that language.
Language Data Source(s) Events/Topics Focused
Amharic (amh) Facebook, X (Twitter) The Tigray War
Arabic (arb)
Existing datasets, Facebook,
News, Reddit, Threads,X (Twitter)
Social issues regading politics and religion
Bengali (ben) YouTube comments Social discourse, generic contemporary topics
Burmese (mya) Existing datasets Kyaw et al. (2024), Wikipedia
Social issues regading politics, ethnicity
and popular culture
Chinese (zho) Tieba, Weibo, Zhihu
Social issues regarding racism,
sexuality/gender and religious discrimination
English (eng) Bluesky, Local news, X (Twitter) US elections and international conflicts
German (deu) Bluesky, Reddit, X (Twitter) COVID-19, and contemporary social issues
Hausa (hau) Facebook, X (Twitter)
Social issues regading politics, ethnicity
and religion
Hindi (hin) Bluesky, Reddit, X (Twitter)
Social issues regarding politics, religion,
and caste
Italian (ita) YouTube, X (Twitter)
Pride parade, immigration crisis
, crime news, Italian justice reform
Khmer (khm)
Facebook, Specialised websites,
Wikipedia, YouTube, Local news
COVID-19, and contemporary social issues
Nepali (nep) Facebook, Local news, X (Twitter) Social discourse, generic contemporary topics
Odia (ori) Bluesky, Local news, X (Twitter) Social discourse, generic contemporary topics
Persian (fas) Bluesky, X (Twitter) Social discourse, generic contemporary topics
Polish (pol) Bluesky, Existing dataset kołos2024banpl COVID-19, and contemporary social issues
Punjabi (pan)
Existing dataset (pnbTenTen)
https://www.sketchengine.eu/pnbtenten-western-punjabi-corpus ,
Youtube Comments,
X (Twitter)
Social issues regarding politics, religion,
and caste
Russian (rus)
Bluesky, X (Twitter),
COVID-19 chatter dataset Banda et al. (2021)
Social discourse, generic contemporary topics
Spanish (spa) Bluesky, X (Twitter)
2010’s inmigration movement,
"Salvemos las dos vidas" movement,
social issues regarding politics
and gender inequality
Swahili (swa) X (Twitter) Kenyan elections
Telugu (tel) Facebook, Reddit, X (Twitter) Social discourse, generic contemporary topics
Turkish (tur) X (Twitter), Existing dataset Çöltekin (2020) Social discourse, generic contemporary topics
Urdu (urd) X (Twitter) Social discourse, generic contemporary topics
Table 7: Summary of dataset composition regarding data sources and events or topics focused.
Appendix C SLMs and LLMs Used
C.1 Multilingual Encoders
C.2 LLMs
Appendix D Experiment Settings
D.1 Experiments settings
For SLMs, we performed language-specific fine-tuning for 3 epochs using a learning rate of 2e-4.
For LLMs in the few-shot setting , we use three shots. ALL prompts are written in English, while the in-context examples are provided in the target language. The full prompts is reported in Appendix F.
All experiments were conducted on local GPU and implemented with EncouRAGe (Strich et al., 2025) framework to execute experiments and measure results.
Appendix E Annotation Guidelines